# Zero and Nonzero Energy Majorana Modes in an Extended Kitaev Chain

### v1

@@ Abstract
This paper studies an extended Kitaev chain with three sublattices per unit cell. This extended version is obtained by hybridizing a modified Su-Schrieffer-Heeger model featuring trimerized unit cells with the standard Kitaev chain, resulting in a hexamer structure on the Majorana basis. Due to the interplay between the sublattice configuration and the $p$-wave superconducting pairing, a rich structure of edge modes beyond the expected Majorana zero modes is obtained. The various Majorana edge modes are further found to demonstrate considerable robustness against some generic perturbations and disorder. The presence of robust Majorana edge modes beyond their zero energy variations potentially offers a step forward in the ongoing efforts to unambiguously detect Majorana modes in experiments.
@@ I Introduction
Motivated by theoretical proposals [11, 12, 13, 14], several experimental studies have reported some signatures of MZMs [15, 16, 17, 18, 19, 20, 21], though such signatures could also be attributed to other mechanisms [22, 23, 24, 25, 26]. At present, unambiguously confirming MZMs in a topological superconductor remains an experimental challenge. Nevertheless, theoretical studies of MZMs and variations of topological superconductors that host them are still actively conducted. For example, despite being Hermitian in nature, MZMs are also found to emerge in a range of non-Hermitian systems [27, 28, 29, 30, 30]. Moreover, periodically driven systems offer a richer playground that enables a different type of Majorana edge modes to emerge in addition to MZMs [31, 32, 33, 34, 35, 36]. It is expected that theoretically exploring these variations of topological superconductors may eventually lead to models whose MZMs could be more unequivocally detected experiments, which therefore serves as an important motivation for such studies.
@@
The SSH model is one of the simplest toy models that supports a topological phase [37]. It describes a 1D chain with nearest-neighbor coupling that alternates in amplitudes between two different values, resulting in the dimerization of the lattice structure (two sites per unit cell). While mathematically very similar to the Kitaev chain, the edge modes in the SSH model are not Majorana fermionic in nature and are thus not as useful for quantum computing purposes. Nevertheless, its simplicity has led to various extensions of the SSH model that involve longer-range coupling [46, 47] and/or a larger unit cell [48, 49, 50, 51], which often result in richer topological features. For example, an extended SSH model with three sites per unit cell supports nonzero energy edge states that are symmetrically placed about zero energy [49, 50].
Drawing inspiration from the above studies of extended SSH models, efforts have been devoted to extending the Kitaev chain as well, leading to topological superconducting models with long-range hopping/pairing [52, 53, 54] and modulated structures such as dimer chains [55, 56]. Our work presents another extension to the Kitaev chain by modulating all system parameters in space with a period of three lattice sites. In the Majorana basis, this results in a lattice structure that exhibits six sites per unit cell. In our proposed model, the system parameters’ modulation is mathematically devised as a tensor product between the regular Kitaev chain [1] and the trimerized SSH chain we recently proposed [50]. By construction, the model exhibits features of long-range hopping and pairing while maintaining a relatively simple mathematical structure. In addition, it preserves fundamental symmetries, including chiral, particle-hole, and time-reversal symmetries, which are all important for its characteristics. This tapestry of particle-hole symmetry and superconducting pairing enriches the system with Majorana edge modes beyond the zero energy excitation that arises alone in the Kitaev chain and its existing extensions [52, 53, 54, 55, 56]. Moreover, the robustness of these Majorana edge modes is highlighted by their persistence under representative generic perturbations we consider below. Remarkably, we find that some of these perturbations actually enrich the system further in terms of its edge states, and therefore its topology.
@@ II.1 Hamiltonian
$\quad\left.+\gamma_{B_{1},j}\gamma_{C_{2},j}\right),$
@@
$\quad\left.+\gamma_{B_{1},j+1}\gamma_{C_{2},j}\right),$
@@
$\quad\left.-\gamma_{C_{2},j}\gamma_{B_{2},j}\right),$
@@
$\quad\left.-\gamma_{B_{2},j+1}\gamma_{C_{2},j}\right),$
@@
II.2 Momentum-space analysis
@@
$\quad+((\delta-\Delta\cos(k))S_{y}+\Delta\sin(k)S_{x})\tau_{x},$ (5)
@@
$S_{x}=\left(\begin{smallmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{smallmatrix}\right),\quad S_{y}=\left(\begin{smallmatrix}0&-i&0\\ i&0&-i\\ 0&i&0\end{smallmatrix}\right).$
@@
$c_{\xi,k}$ $=\sum_{j}e^{-ikj}c_{\xi,j},\quad c_{\xi,k}^{\dagger}=\sum_{j}e^{ikj}c_{\xi,j}^{\dagger},$
$c_{\xi,-k}$ $=\sum_{j}e^{ikj}c_{\xi,j},\quad c_{\xi,-k}^{\dagger}=\sum_{j}e^{-ikj}c_{\xi,j}^{\dagger}.$
@@
$\mu$ $=$ $\sqrt{2\left((J_{1}+J_{2})^{2}-(\delta-\Delta)^{2}\right)},\quad\text{at }k=0,$
$\mu$ $=$ $\sqrt{2\left((J_{1}-J_{2})^{2}-(\delta+\Delta)^{2}\right)},\quad\text{at }k=\pi.$ (8)
Figure 2 shows some representative energy spectra in momentum space. When the system parameters satisfy Eq. (8), band touching points at $k=0$ and $k=\pm\pi$, respectively, are observed (see Figs. 2(a) and 2(b)). In contrast, at arbitrary parameter values, the system exhibits a fully gapped spectrum (see Fig. 2(c)).
@@
II.3 Real-space analysis and topological invariant
@@
That both types of edge modes coexist in our system is a consequence of our construction that mathematically combines the two models. It is also worth noting that, unlike the extended SSH models of Ref. [50], the nonzero edge modes in our system are also Majorana in nature like their zero energy counterparts. Therefore, such nonzero edge modes may potentially contribute to quantum computing applications or, at the very least, supplement the current experimental efforts in unambiguously detecting Majorana modes.
Figure 3: (a) and (c) are the energy spectrum as a function of $\mu$ and $\Delta$, respectively, for $N=60$ unit cells. (b) and (d) show the numerically calculated winding number associated with (a) and (c), respectively. (a) and (b) are evaluated at $J_{1}=0.5$, $J_{2}=2.5$, and $\delta=\Delta=0.4$. (c) and (d) are evaluated at $J_{1}=0.5$, $J_{2}=2.5$, $\delta=0.4$, and $\mu=2$. The black vertical lines mark the locations of the band touching points anticipated by Eq. (8). The red vertical dashed line indicates the value used in Fig. 4.
@@
Intuitively, Eq. (11) is defined such that each combination of the sublattice, particle-hole, and lattice site degrees of freedom is assigned a unique integer, thereby allowing us to obtain the complete components of each edge mode. Figure 4 reveals that while all Majorana edge modes are highly localized, the nonzero energy edge modes only have two dominating peaks near a system’s edge (see Fig. 4(a,b,d,e)). By contrast, MZMs are found to have significant support on more degrees of freedom near a system’s edge. The considerably distinct features of the different types of edge modes in principle enable their separate experimental detection.
Figure 4: Wave function profiles associated with the system’s edge states for $N=60$ unit cells at $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, and $\mu=3.5$. Note that each subplot displays one left-localized edge state and one right-localized edge state.
@@ IV.1 Effects of perturbation
Figures 5 and 6 summarize our findings by respectively plotting the energy spectrum in momentum space and real space in the presence of the aforementioned perturbations. In the momentum space spectrum of Fig. 5, the expected energies of the edge modes when OBC are applied are also plotted to highlight the spectral separation of the edge states from the bulk bands. Meanwhile, the real space spectrum of Fig. 6 explicitly highlights the fate of the system’s edge states as the perturbation strengths are varied. Together, both figures not only demonstrate the topological robustness of our system, but also highlight specific features that emerge as a result of the applied perturbations.
@@
Figure 6: Energy spectrum with respect to all five perturbation strengths for $N=60$ unit cells. The parameter values for all five cases are $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, and $\mu=3.5$. The red vertical dashed lines mark the parameter values used in Fig. 5.
@@ IV.2 Effects of disorder
Figure 7: Energy spectrum with respect to the disorder strength for $N=60$ unit cells. The starting parameter values for all five cases are $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$ and $\mu=3.5$. (a) Disorder acts on $J_{1}$, (b) Disorder acts on $J_{2}$, (c) Disorder acts on $\delta$, (d) Disorder acts on $\Delta$, and (e) Disorder acts on $\mu$. Each data point is averaged over 100 disorder realizations.
Figure 8: Energy spectrum with respect to the disorder strength for $N=60$ unit cells. The starting parameter values for all five cases are $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, $\mu=3.5$, and $\alpha_{o}$ = $\beta_{o}$ = $\zeta_{o}$= $\vartheta_{o}$ = $\varrho_{o}$ = 0.5. (a) Disorder acts on $\alpha_{o}$, (b) Disorder acts on $\beta_{o}$, (c) Disorder acts on $\zeta_{o}$, (d) Disorder acts on $\vartheta_{o}$, and (e) Disorder acts on $\varrho_{o}$. Each data point is averaged over 100 disorder realizations.

### v2

@@ Abstract
This paper studies an extended Kitaev chain with three sublattices per unit cell. This extended version is obtained by hybridizing a modified Su-Schrieffer-Heeger model featuring trimerized unit cells with the standard Kitaev chain, resulting in a hexamer structure on the Majorana basis. Due to the interplay between the sublattice configuration and the $p$-wave superconducting pairing, a rich structure of edge modes beyond the expected Majorana zero modes is obtained. The various Majorana edge modes are further found to demonstrate considerable robustness against some generic perturbations and disorder. The presence of the non-zero Majorana edge modes in our system has the potential advantage that they could, in principle, be more unambiguously detected as compared to their zero energy counterparts, the detection of which remains an open problem. Therefore, while our system does not directly solve this open problem, it potentially offers a route to an intermediate solution that involves unambiguously detecting non-zero energy Majorana edge modes instead.
@@ I Introduction
Motivated by theoretical proposals [59, 50, 44, 56], several experimental studies have reported some signatures of MZMs [19, 54, 22, 16, 57, 51, 49], though such signatures could also be attributed to other mechanisms [65, 53, 60, 47, 43, 75]. At present, unambiguously confirming MZMs in a topological superconductor remains an experimental challenge. Nevertheless, theoretical studies of MZMs and variations of topological superconductors that host them are still actively conducted. For example, despite being Hermitian in nature, MZMs are also found to emerge in a range of non-Hermitian systems [79, 27, 46, 83]. Moreover, periodically driven systems offer a richer playground that enables a different type of Majorana edge modes to emerge in addition to MZMs [35, 73, 84, 77, 5, 11]. It is expected that theoretically exploring these variations of topological superconductors may eventually lead to models whose MZMs could be more unequivocally detected experiments, which therefore serves as an important motivation for such studies.
@@
The SSH model is one of the simplest toy models that supports a topological phase [69]. It describes a 1D chain with nearest-neighbor coupling that alternates in amplitudes between two different values, resulting in the dimerization of the lattice structure (two sites per unit cell). While mathematically very similar to the Kitaev chain, the edge modes in the SSH model are not Majorana fermionic in nature and are thus not as useful for quantum computing purposes. Nevertheless, its simplicity has led to various extensions of the SSH model that involve longer-range coupling [29, 23, 17] and/or a larger unit cell [39, 4, 30, 31], which often result in richer topological features. For example, an extended SSH model with three sites per unit cell supports nonzero energy edge states that are symmetrically placed about zero energy [4, 30].
Drawing inspiration from the above studies of extended SSH models, efforts have been devoted to extending the Kitaev chain as well, leading to topological superconducting models with long-range hopping/pairing [74, 2, 62, 52, 18] and modulated structures such as dimer chains [76, 78]. Our work presents another extension to the Kitaev chain by modulating all system parameters in space with a period of three lattice sites. In the Majorana basis, this results in a lattice structure that exhibits six sites per unit cell. In our proposed model, the system parameters’ modulation is mathematically devised as a tensor product between the regular Kitaev chain [36] and the trimerized SSH chain we recently proposed [30]. By construction, the model exhibits features of long-range hopping and pairing while maintaining a relatively simple mathematical structure. In addition, it preserves fundamental symmetries, including chiral, particle-hole, and time-reversal symmetries, which are all important for its characteristics. This tapestry of particle-hole symmetry and superconducting pairing enriches the system with Majorana edge modes beyond the zero energy excitation that arises alone in the Kitaev chain and its existing extensions [74, 2, 62, 52, 18, 76, 78]. Moreover, the robustness of these Majorana edge modes is highlighted by their persistence under representative generic perturbations we consider below. Remarkably, we find that some of these perturbations actually enrich the system further in terms of its edge states, and therefore its topology.
@@ II.1 Hamiltonian
$\qquad\left.+\gamma_{B_{1},j}\gamma_{C_{2},j}\right),$
@@
$\qquad\left.+\gamma_{B_{1},j+1}\gamma_{C_{2},j}\right),$
@@
$\qquad\left.-\gamma_{C_{2},j}\gamma_{B_{2},j}\right),$
@@
$\qquad\left.-\gamma_{B_{2},j+1}\gamma_{C_{2},j}\right),$
@@
To intuitively demonstrate the presence of Majorana modes in our system, we first consider the special parameter values $\mu=J_{1}=\delta=0$. As schematically depicted in Fig. 2(a), a pair of Majorana fermions at each edge are decoupled from each other and from the rest of the chain, which thus correspond to MZMs. However, by tuning $\mu$ to a finite value while still keeping $J_{1}=\delta=0$ (see Fig. 2(b)), the two disconnected MZMs at each edge are now coupled with each other, though they are still decoupled from the rest of the chain. In this case, the linear combination of the two leftmost Majorana operators, i.e., $\gamma_{L}=a\gamma_{A_{1},1}+b\gamma_{A_{2},1}$, is closed under commutator with the system’s Hamiltonian. The equation $H,\gamma_{L}]=E\gamma_{L}$ can then be rewritten as an eigenvalue equation involving a $2\times 2$ matrix, the two solutions of which correspond to non-zero energy edge modes. Therefore, like the regular Kitaev chain, special parameter values exist in which the Majorana edge modes are easily tractable. However, while the regular Kitaev chain only supports MZMs, our model is expected to support non-zero energy edge modes, as we also numerically confirm later on.
@@
Figure 2: (a) Schematic diagram of the first and last unit cells in a chain of N=100 unit cells at $\mu=J_{1}=\delta=0$. Note that there are two Majorana modes decoupled at each end of the chain. (b) Schematic diagram of the first and last unit cells in a chain of N=100 unit cells at $J_{1}=\delta=0$. Note that there are two Majorana modes coupled at each end of the chain, but they are disconnected from the rest of the chain.
II.2 Momentum-space analysis and topological invariant
@@
$\qquad+((\delta-\Delta\cos(k))S_{y}+\Delta\sin(k)S_{x})\tau_{x},$ (5)
@@
$S_{x}=\left(\begin{smallmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{smallmatrix}\right),\qquad S_{y}=\left(\begin{smallmatrix}0&-i&0\\ i&0&-i\\ 0&i&0\end{smallmatrix}\right).$
@@
$c_{\xi,k}$ $=\sum_{j}e^{-ikj}c_{\xi,j},\qquad c_{\xi,k}^{\dagger}=\sum_{j}e^{ikj}c_{\xi,j}^{\dagger},$
$c_{\xi,-k}$ $=\sum_{j}e^{ikj}c_{\xi,j},\qquad c_{\xi,-k}^{\dagger}=\sum_{j}e^{-ikj}c_{\xi,j}^{\dagger}.$
@@
$\mu$ $=$ $\sqrt{2\left((J_{1}+J_{2})^{2}-(\delta-\Delta)^{2}\right)},\qquad\text{at }k=0,$
$\mu$ $=$ $\sqrt{2\left((J_{1}-J_{2})^{2}-(\delta+\Delta)^{2}\right)},\qquad\text{at }k=\pi.$ (8)
Figure 3 shows some representative energy spectra in momentum space. When the system parameters satisfy Eq. (8), band touching points at $k=0$ and $k=\pm\pi$, respectively, are observed (see Figs. 3(a) and 3(b)). In contrast, at arbitrary parameter values, the system exhibits a fully gapped spectrum (see Fig. 3(c)).
@@
As the preceding discussion only presents an invariant that captures the presence/absence of the MZMs, one may wonder if the non-zero energy Majorana edge modes can be characterized by a similar invariant. Unfortunately, developing such an invariant is a highly nontrivial task. In principle, a suitable topological invariant for the non-zero energy Majorana edge modes could be systematically constructed via some transfer matrix method, such as in the spirit of Refs. [20, 41]. In these studies, the transfer matrix method was applied to identify MZMs by solving some matrix equation that yields localized, zero energy solutions [20, 41]. However, adapting the same method to identifying the non-zero energy edge modes in our model is significantly more challenging due to the $6\times 6$ matrix involved here as compared with the $2\times 2$ matrix in previous studies. Moreover, most importantly, the non-zero energy edge modes lack an exact, closed-form expression for their energy within the bulk gap, which is required to extract the conventional transfer matrix. These aspects render any analytical calculation extremely difficult, if not impossible, to carry out. Therefore, the topological nature of the non-zero energy edge modes will be established through another means, i.e., by directly computing their spatial profiles in the following section and comparing them from the expected profiles at special parameter values.
II.3 Real-space analysis
@@
That both types of edge modes coexist in our system is a consequence of our construction that mathematically combines the two models. It is also worth noting that, unlike the extended SSH model of Ref. [30], the nonzero edge modes in our system are also Majorana in nature like their zero energy counterparts. Therefore, such nonzero edge modes may potentially contribute to quantum computing applications or, at the very least, supplement the current experimental efforts in unambiguously detecting Majorana modes.
A fundamental characteristic of the Kitaev chain is the emergence of MZMs within the topological phase, which exists for chemical potentials $|\mu|<2|t|$, where $t$ is the hopping parameter. A topological phase transition occurs precisely at $|\mu|=2|t|$, where the pairing parameter does not affect the transition point, as long as it is nonzero. Since there is only a single phase transition point, the MZMs emerge at starting point of $\mu=0$, marking the boundary of this phase, i.e., from $\mu=0$ to $|\mu|=2|t|$. Unlike the regular Kitaev chain, the extended model presented in this paper exhibits a richer phase diagram, characterized by two distinct topological phase transition points, as defined by Eq. (8). This structure introduces a critical advantage. That is, the region supporting MZMs is not fixed to a specific range of $\mu$. Instead, both of the phase transition points can be shifted to different ranges of $\mu$ and tuned by adjusting the system’s coupling strengths, as the phase transition points depend on all system’s parameters (see Eq. (8)). This feature is absent in the regular Kitaev chain, where the onset of MZMs is fixed at $\mu=0$.
The primary experimental signature for MZMs is a zero-bias conductance peak in tunneling spectroscopy [55]. At the zero-bias peak, trivial excitations such as the Andreev-bound state [47, 43, 75] and the Kondo effect [40, 45] can arise. Such trivial excitations induce zero-bias anomalies that strongly resemble the tunneling signatures often associated with MZMs, posing an experimental challenge to their identification. Such trivial excitations are also expected to take place at finite energy. Likewise, their presence may hamper the detection of the non-zero edge modes. The non-zero edge modes in our system appear as finite-bias peaks in tunneling spectroscopic measurements, and the characteristics of these peaks depend on the system’s parameters. Unlike MZMs, which are uniquely identified by a zero-bias peak, the finite-energy Majorana modes provide experimental flexibility by means of their pair of peaks, in principle, that occur at energy $\pm E$ within the superconducting gap. A key strategy is to take advantage of the system’s parameter dependency. That is, varying the system parameters will shift or eliminate peaks of trivial origin, while the pair of peaks from finite-energy Majorana modes should evolve coherently and persist across the range of parameters chosen. This enables tracking the non-zero Majorana modes through several experiments across numerous parameter regimes. It is expected that this means of unambiguously detecting non-zero-energy Majorana modes provides an alternative, potentially more practical, pathway to exploring Majorana physics before MZMs unambiguous detection is possible.
Figure 4: (a) and (c) are the energy spectrum as a function of $\mu$ and $\Delta$, respectively, for $N=100$ unit cells. (b) and (d) show the numerically calculated winding number associated with (a) and (c), respectively. (a) and (b) are evaluated at $J_{1}=0.5$, $J_{2}=2.5$, and $\delta=\Delta=0.4$. (c) and (d) are evaluated at $J_{1}=0.5$, $J_{2}=2.5$, $\delta=0.4$, and $\mu=2$. (e) and (f) show zoomed-in views of the MZMs from (a) and (c), respectively. The black vertical lines mark the locations of the band touching points anticipated by Eq. (8). The red vertical dashed line indicates the value used in Fig. 5.
@@
Intuitively, Eq. (11) is defined such that each combination of the sublattice, particle-hole, and lattice site degrees of freedom is assigned a unique integer, thereby allowing us to obtain the complete components of each edge mode. Figure 5 reveals that while all Majorana edge modes are highly localized, the nonzero energy edge modes only have two dominating peaks near a system’s edge (see Fig. 5(a,b,d,e)). This feature is consistent with the intuitive picture of the model at special parameter values $J_{1}=\delta=0$ previously elucidated in Sec. II.1, where the first and last two Majorana operators hybridize with each other (while still decoupled from the rest of the chain) and form non-zero energy Majorana edge modes. That the significant two peaks are also observed away from the special parameter values highlights the topological nature of the non-zero energy edge modes. By contrast, MZMs are found to have significant support on more degrees of freedom near a system’s edge. The considerably distinct features of the different types of edge modes in principle enable their separate experimental detection. To further ensure that these characteristics do not depend on the number of lattice sites in the system, we show the wave function profiles of the edge modes at different system sizes in Appendix A.
Figure 5: Wave function profiles associated with the system’s edge states for $N=100$ unit cells at $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, and $\mu=3.5$. Note that each subplot displays one left-localized edge state and one right-localized edge state.
@@ IV.1 Effects of perturbation
Figures 6 and 7 summarize our findings by respectively plotting the energy spectrum in momentum space and real space in the presence of the aforementioned perturbations. In the momentum space spectrum of Fig. 6, the expected energies of the edge modes when OBC are applied are also plotted to highlight the spectral separation of the edge states from the bulk bands. Meanwhile, the real space spectrum of Fig. 7 explicitly highlights the fate of the system’s edge states as the perturbation strengths are varied. Together, both figures not only demonstrate the topological robustness of our system, but also highlight specific features that emerge as a result of the applied perturbations.
@@
Figure 7: Energy spectrum with respect to all five perturbation strengths for $N=100$ unit cells. The parameter values for all five cases are $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, and $\mu=3.5$. The red vertical dashed lines mark the parameter values used in Fig. 6.
@@ IV.2 Effects of disorder
Figure 8: Energy spectrum with respect to the disorder strength for $N=100$ unit cells. The starting parameter values for all five cases are $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$ and $\mu=3.5$. (a) Disorder acts on $J_{1}$, (b) Disorder acts on $J_{2}$, (c) Disorder acts on $\delta$, (d) Disorder acts on $\Delta$, and (e) Disorder acts on $\mu$. Each data point is averaged over 100 disorder realizations.
Figure 9: Energy spectrum with respect to the disorder strength for $N=100$ unit cells. The starting parameter values for all five cases are $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, $\mu=3.5$, and $\alpha_{o}$ = $\beta_{o}$ = $\zeta_{o}$= $\vartheta_{o}$ = $\varrho_{o}$ = 0.5. (a) Disorder acts on $\alpha_{o}$, (b) Disorder acts on $\beta_{o}$, (c) Disorder acts on $\zeta_{o}$, (d) Disorder acts on $\vartheta_{o}$, and (e) Disorder acts on $\varrho_{o}$. Each data point is averaged over 100 disorder realizations.